# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

### v1

@@ Abstract
While lattice Boltzmann methods were originally derived as the Boltzmann limit of a Boolean lattice gas, modern improvements to lattice Boltzmann methods have broken this link. Because of this most lattice Boltzmann methods are not unconditionally stable and extending them to model fluctuations is challenging. In this article we show that the integer lattice gas method developed by Blommel and Wagner has a lattice Boltzmann limit corresponding to entropic lattice Boltzmann methods, while also recovering proper ideal gas equilibrium fluctuations. We also show that it is possible to accelerate these integer lattice gases such that they become computationally competitive with the corresponding lattice Boltzmann method.
@@ I Introduction
Frisch, Hasslacher and Pomeau’s Frisch et al. (1987) extension of lattice gas automata to allow for the simulation of hydrodynamics opened up the new field of discrete simulations of fluid dynamics. Despite their very simple local evolution rules these lattice gases could be shown to recover the continuity and Navier-Stokes equations. The fluctuations of these lattice gases closely resembled the fluctuations of real liquids Rivet and Boon (2001).
@@
Initially the Boltzmann limit was applied to lattice gas systems to analyze their hydrodynamic limit in analogy to the procedure in kinetic theory Huang (1987). This Boltzmann limit consists of taking an ensemble average that removes the fluctuations. McNamara et al. McNamara and Zanetti (1988) were the first to use this Boltzmann limit of the lattice gas as a simulation method in its own right. Because simulating the lattice Boltzmann equation eliminated statistical noise, this was an advantage for simulating noise-free hydrodynamics. Previously this had to be achieved by performing a statistical average over several lattice gas simulations.
@@
The lattice Boltzmann method was later decoupled from the lattice gas entirely when Qian, d’Humieres, and Lallemand Qian et al. (1992) proposed a simplified BGK collision operator. By simply choosing a local equilibrium distribution (the now ubiquitous second order polynomial equilibrium distribution) they simplified the algorithm and made it significantly more accessible. They were also able to remove some non-Galilean invariant artifacts of Boolean lattice gas systems, such as a velocity dependent viscosity Frisch et al. (1987); Higuera and Jimenez (1989) and pressure, even if these deviations became most notable in parameter ranges (i.e. at high Mach numbers) that were still not accessible to these lattice Boltzmann methods.
@@
The simplicity of this BGK collision operator, more than anything else, influenced the explosive rise of lattice Boltzmann methods over their ancestral lattice gas methods. Certain systems, however, such as the Brownian motion of colloids considered by Ladd et al. Ladd et al. (1988), require fluctuations in order to be simulated. These fluctuating systems were, at the time, limited to being simulated by lattice gases. To allow these simulations to benefit from the advances of lattice Boltzmann methods, Ladd developed a fluctuating lattice Boltzmann method. The basic idea was to implement Landau and Lifshitz’ Landau et al. (1959) fluctuating stress tensor into the collision operator leading to a Langevin approach Lemons and Gythiel (1997) for the fluctuating lattice Boltzmann method. This fluctuating lattice Boltzmann method recovered the fluctuating hydrodynamic equations Ladd (1993). However, utilizing the fluctuating stress tensor alone limited the recovery of the correct fluctuations to those with a wavelength on the order of the system size, while the total order of fluctuations was suppressed by about a factor of two.
@@
Adhikari et al. discovered that in order to recover the smaller scale fluctuations, a full set of fluctuating modes, rather than just those associated with hydrodynamic quantities, was needed Adhikari et al. (2005). This method is restricted to ideal gas systems. For non-ideal systems additional corrections are required Belardinelli et al. (2015). Since the Langevin approach is based on continuous densities, the large fluctuations at low densities will lead to the formation of non-physical negative densities (Kähler, 2012, sec. 5.3).
@@
When considering methods for simulating fluctuating systems, however, it is much more natural to think in terms of discrete systems since fluctuations fundamentally arise because of the discreteness of nature Einstein (1906). This inspired Chopard et al. to instead develop novel lattice gas algorithms instead of adding fluctuations artificially to the lattice Boltzmann algorithm (Chopard et al., 1998; Chopard and Droz, 1998, sec. 5.7). They extended the occupation numbers of the lattice gas from Boolean values to integer values, a concept that was first introduced by Chen et al. the year prior with their Digital Physics approach Chen et al. (1997). This allowed for the lattice gas system of Chopard et al. to recover a Poisson distribution in global equilibrium. Secondly, Chopard et al. implemented a sampling algorithm, which was originally theorized by Boghosian et al. Boghosian et al. (1997), that distributed particles randomly across the system velocities based on a local equilibrium ensemble. For the diffusive case, a multinomial distribution over the occupation numbers themselves was utilized for the local equilibrium ensemble, which was sufficient to follow the conservation of mass required by the system (Chopard and Droz, 1998, sec. 5.7).
@@
In extending the method to hydrodynamic systems, however, Chopard et al. did not utilize an explicit local equilibrium ensemble due to the complexity involved in deriving this distribution while accounting for momentum conservation. Instead, a multinomial distribution similar to that of the diffusive case was used, centered around the second order polynomial equilibrium distribution. This on its own would violate momentum conservation, so a random factor that redistributed a single particle at a time until momentum conservation was upheld was also added Chopard et al. (1998). Both of these methods were, for higher densities, well approximated by Gaussian distributions, which caused their algorithmic complexity to be approximately O(1) with respect to the system density, matching the complexity of the lattice Boltzmann algorithm.
@@
Blommel and Wagner created another integer based lattice gas algorithm to simulate hydrodynamic systems Blommel and Wagner (2018). Their algorithm utilized a series of random two-particle collisions in order to redistribute particle velocities while still maintaining momentum conservation with collision rules explicitly derived by imposing detailed balance on the system. This allowed Blommel and Wagner to recover Poisson distributed fluctuations in global equilibrium similar to Chopard et al..
@@
When analyzing their algorithm, the global equilibrium average distribution was found to correspond not to the polynomial lattice Boltzmann equilibrium distribution, but to the entropic lattice Boltzmann equilibrium distribution. This equilibrium distribution had already been derived by Ansumali et al. Ansumali et al. (2003) previously by minimizing an H-functional. The polynomial equilibrium distribution utilized by Chopard et al. Chopard et al. (1998) for their hydrodynamic multiparticle lattice gas, therefore, was actually incompatible with a microscopic lattice gas implementation.
@@
Since the number of individual two-particle collisions needed to achieve a state close to local equilibrium scales with the square of the number of particles at each lattice site, Blommel and Wagner’s algorithm proved much slower than the fluctuating lattice Boltzmann algorithm even for moderate local densities of a few hundred particles per lattice site. This lead Seekins and Wagner Seekins and Wagner (2022) to accelerate Blommel’s algorithm. A sampling method similar to that of Chopard et al. was implemented in order to reduce the number of random numbers needed, leading to increased efficiency.
@@
Using detailed balance in order to derive a local equilibrium distribution, Seekins and Wagner recovered an identical multinomial distribution to that of Chopard et al. (Chopard and Droz, 1998, sec. 5.7). Although this meant that the diffusive algorithm created by Seekins and Wagner was essentially identical to that of Chopard et al., the derivation of the equilibrium distribution via detailed balance was novel, and could be used to derive the proper equilibrium distribution for hydrodynamic systems. The efficiency of the algorithm was comparable to a diffusive fluctuating lattice Boltzmann algorithm, along with having a complexity at high densities of approximately O(1) with respect to density Seekins and Wagner (2022); Wagner and Strand (2016).
@@
In this paper, we extend our prior work for diffusive systems to hydrodynamic systems by deriving an appropriate local equilibrium distribution. for simplicity, we limit the system to one dimension, leaving an extension to higher dimensions to future work. The paper is structured as follows: in Section II of the paper, we will cover the sampling lattice gas algorithm in detail. In Section III we derive the local equilibrium distribution and provide an analysis of the Poisson distribution as a global equilibrium distribution for the integer lattice gas. Section IV will discuss the Boltzmann limit of the lattice gas and its uniqueness. Section V covers the results of two test cases used to study the non-equilibrium dynamics of the system. Finally, section VI will compare the efficiency of our sampling lattice gas to that of both Blommel and Wagner and an entropic lattice Boltzmann algorithm.
@@ II.1 Sampling Lattice Gas Basics
A lattice gas consists of a set of discrete cells situated on a lattice, each containing a discrete number of velocity states $v_{i}$ for particles to occupy Wolf-Gladrow (2004). The velocities are usually identical for every cell, and are defined such that if $x$ is a position on the lattice, then for any velocity $v_{i}$, $x+v_{i}$ is also a position on the lattice, assuming a timestep of $\Delta t=1$. These velocities are also typically restricted in number such that the particle displacement over a single timestep is small. The particles are accounted for by the occupation numbers $n_{i}(x,t)$, denoting the particles that move from lattice site $x-v_{i}$ at time $t-1$ to lattice site $x$ at time $t$. These particles are then redistributed among the occupation numbers, observing the applicable conservation rules, and streamed. This two-step process represents a discrete timestep for the system. Thus, the $n_{i}(x,t)$ evolve according to,
@@
which is known as the lattice gas evolution equation. Here $\{n_{j}\}_{g}$ represents the set of all $n_{i}$ of the system at the current time. In this evolution equation, the collision operator $\Xi_{i}$ is a stochastic operator. The explicit nature of the $n_{i}$ is dependent on the lattice gas system observed. For Boolean lattice gases, for example, these $n_{i}$ would be Boolean in nature, leading to $\Xi_{i}$ to be calculable explicitly within a lookup table due to the limited nature of the probability space for single-occupation velocity states Frisch et al. (1986, 1987). In our integer lattice gas method, the $n_{i}$-values become integer based, which leads to a collision operator $\Xi_{i}$ that can take on integer values.
In the ideal gas systems considered here the collision operator will only act on the local distribution of occupation numbers, $\{n_{j}\}_{l}$, which consists only of the local occupation numbers at position $x$ and $t$. In this nomenclature we exclude the implicit dependence on $x$ and $t$ to make the equations more readable. However, in principle the collision operator can depend on the global distribution $\{n_{j}\}_{g}$, as would be the case for non-ideal lattice gas systems. An example of that would be Rothmann and Keller’s algorithm Rothman and Zaleski (1994) that is now known as the color gradient method. This method depends on neighboring site average density gradients.
The sampling collision operator takes inspiration from the BGK collision operator for the Boltzmann equation. The Boltzmann equation is an evolution equation for a distribution function $f_{i}(x,t)$, and the BGK collision operator is an approximation that simply moves the system towards it’s local equilibrium distribution $f_{i}^{0}(x,t)$ Bhatnagar et al. (1954) where the local equilibrium distribution depends on the locally conserved quantities. This BGK lattice Boltzmann algorithm will be discussed in more detail in upcoming sections.
In a lattice gas system, the system is described by $\{n_{i}\}_{g}$. Unlike in the lattice Boltzmann case, the local equilibrium state for the $n_{i}$ is not given by an equilibrium distribution, but rather by a probability distribution around some average value, known as the local equilibrium ensemble. This local equilibrium ensemble, given a local pre-collision set of occupation numbers $\{n_{i}\}_{l}$, gives a probability $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ for each post-collision set $\{\hat{n}_{i}\}_{l}$. A post-collision set of occupation numbers can be then be sampled from the local equilibrium ensemble in a process discussed in detail in Seekins and Wagner’s paper Seekins and Wagner (2022). The collision operator for the sampling lattice gas algorithm thus becomes the difference between the pre-collision and post-collision occupation numbers, explicitly defined as,
@@
However, this representation of the collision operator does not account for the fact that the system may not relax fully to local equilibrium in one timestep. In order to simulate this under-relaxation in a system of discrete particles, we can represent the fraction of particles that should undergo collision in that timestep as $\omega$. We create a subset of our occupation numbers $\{n^{\omega}_{i}\}_{l}$ to undergo collision, with the probability that any single particle is a member of the subset being given by $\omega$. Therefore, each specific value of $n_{i}^{\omega}$ has a probability associated with it given by the binomial distribution,
@@
It remains to derive the local equilibrium ensemble explicitly.
@@ II.2 Derivation of the Local Equilibrium Ensemble
In the hydrodynamic case considered here there are two conserved quantities: mass and momentum. Since momentum is conserved, the probability distribution for the $n_{i}$ will be correlated. However, by transforming our representation of the system from being expressed as an set of individual occupation numbers $\{n_{i}\}_{l}$ into being expressed by the hydrodynamic moments we may be able to obtain an uncorrelated probability distribution. These moments are physical quantities within the simulation, with a total number given by the total number of system velocities $Q$. From some fundamental particle collision process like that of Blommel and Wagner Blommel and Wagner (2018) we deduce the local equilibrium probability distributions for the non-conserved moments. We then use these local equilibrium ensembles to sample the post-collision moments as random variables. This is the extent of how far we are able to describe our algorithm in all generality.
In order to know what moment representation is applicable to our system, we must focus our attention on a specific example. As this paper serves as a proof of concept for our sampling algorithm, we focus on the simplest possible system that recovers hydrodynamic effects, a one-dimensional system with three total velocity vectors, leaving the extension of the algorithm to more complex systems, including higher dimensional systems, to future work.
@@
Thus, we can redefine $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ in terms of $N$, $J$, and $\pi$ as $P^{0}(\hat{N},\hat{J},\hat{\pi};N,J,\pi)$, or the probability of finding a system in a state with moments $\hat{N}$, $\hat{J}$, and $\hat{\pi}$ after a collision given a pre-collision state with moments $N$, $J$, and $\pi$. We can then simplify this representation, as our conservation rules state that $\hat{N}=N$ and $\hat{J}=J$. Furthermore, because we are assuming that the system is in local equilibrium when described by this distribution, we know that $\hat{\pi}$ does not depend on the pre-collision $\pi$-value. The probability distribution for our ensemble can then be written as $P^{0}(\pi;N,J)$, which is the equilibrium probability of finding the system in a post-collision state with a $\pi$-value of $\pi$, given a pre-collision system with conserved moments $N$ and $J$, and an arbitrary initial value for $\pi$.
We now derive the particular probability distribution that recovers the integer lattice gas results obtained by Blommel and Wagner. In local equilibrium, this distribution must be invariant under collisions. By using Blommel and Wagner’s algorithm as a basis, we restrict the possible collisions of our system to only involving pairs of particles. These collisions conserve momentum, and their equilibrium ensemble must be stationary over each collision. Thus, we can look at the equilibrium distribution that corresponds to Blommel and Wagner’s algorithm as the effect of many individual two-particle collisions.
Due to $P^{0}(\pi;N,J)$ being invariant under a single collision we can write,
$T_{\pi\to\pi+2}P^{0}(\pi;N,J)=T_{\pi+2\to\pi}P^{0}(\pi+2;N,J),$ (12)
where $T_{\pi\to\pi+2}$ and $T_{\pi+2\to\pi}$ are the transition probabilities between states $\pi$ and $\pi+2$ under a single two-particle collision. These transition probabilities represent the probability of any collision occurring that increases or decreases the local $\pi$-value respectively. In order to write these transition probabilities explicitly, we take the sum of the probabilities of each possible collision that increases or decreases the $\pi$-value depending on which transition probability we wish to write. There are only four two-particle collisions for a D1Q3 system that alter the $\pi$-value, two that increase the $\pi$-value by two and two that decrease it by two,
$\pi\rightarrow\pi+2\biggr{\{}\begin{array}[]{c}\{0,0\}\to\{1,-1\}\\ \{0,0\}\to\{-1,1\},\\ \end{array}\hskip 5.69046pt$ (13)
$\pi+2\rightarrow\pi\biggr{\{}\begin{array}[]{c}\{-1,1\}\to\{0,0\}\\ \{1,-1\}\to\{0,0\}.\\ \end{array}\hskip 8.5359pt$ (14)
The probabilities of each of these possible collisions occurring is given by the probability of picking a pair of particles with the proper set of velocity indices $\{i,j\}$, which is written as $p_{\{i,j\}}$, multiplied by the probability of the collision occurring from velocity state $\{i,j\}$ to velocity state $\{k,l\}$, written as $P_{\{i,j\}\to\{k,l\}}$. Thus, $T_{\pi\to\pi+2}$ and $T_{\pi+2\to\pi}$ can be defined as,
@@
where $\delta_{ij}$ is the Kronecker delta, and is incorporated to avoid double counting particles should the indices $i$ and $j$ be the same. The $P_{\{i,j\}\to\{k,l\}}$-values were defined explicitly by Blommel and Wagner such that at zero velocity the typical lattice Boltzmann equilibrium values are recovered on average Blommel and Wagner (2018). For the relevant collisions in Eqs. (15) and (16), these values are given by,
@@
It is not obvious that the probability $P_{\{i,j\}\to\{k,l\}}$ should necessarily be independent of the locally conserved quantities. However, as Blommel and Wagner’s method makes this assumption, we make it as well in order to adapt their algorithm. A true test of the validity of this approach could be done by looking at the coarse-graining of a molecular dynamics simulation, similar to the work done by Parsa et al. Parsa and Wagner (2017).
@@
Eq. (12) can then be rewritten as,
@@
fully describes the local equilibrium ensemble. If we sample random $\hat{\pi}$-values out of this local equilibrium ensemble, our collision operator in moment space becomes,

### v2

@@ Abstract
This paper constitutes a step in the direction of developing integer lattice gas methods as an attractive alternative to lattice Boltzmann methods. Here we show that to Boltzmann limit the one dimensional Blommel integer lattice gas is very close to entropic lattice Boltzmann. More interestingly the integer lattice gas retains additional correlations that prevent the existence of a well defined Boltzmann limit. In the analysis of the decaying sine wave we will see that in some situations the bulk viscosity can crucially depend on such correlations beyond the Boltzmann limit. A sampling collision operator, introduced here, can speed up the execution time to make the algorithm obtain comparable computational efficiency to entropic lattice Boltzmann methods.
@@ I Introduction
Lattice gas methods for simulating hydrodynamics have a long history, but they have fallen out of favor compared to lattice Boltzmann methods which were originally derived from lattice gas methods. We will give a brief history of this development in our introduction, and then argue that novel integer lattice gas approaches may both be competitive with lattice Boltzmann methods and allow for additional correlations and fluctuations lost in the Boltzmann description to be retained.
@@
Frisch, Hasslacher and Pomeau’s [1] extension of lattice gas automata to allow for the simulation of hydrodynamics opened up the new field of discrete simulations of fluid dynamics. Despite their very simple local evolution rules these lattice gases could be shown to recover the continuity and Navier-Stokes equations. The fluctuations of these lattice gases closely resembled the fluctuations of real liquids [2].
@@
Initially the Boltzmann limit was applied to lattice gas systems to analyze their hydrodynamic limit in analogy to the procedure in kinetic theory [3]. This Boltzmann limit consists of taking an ensemble average that removes the fluctuations. McNamara et al. [4] were the first to use this Boltzmann limit of the lattice gas as a simulation method in its own right. Because simulating the lattice Boltzmann equation eliminated statistical noise, this was an advantage for simulating noise-free hydrodynamics. Previously this had to be achieved by performing a statistical average over several lattice gas simulations.
@@
The lattice Boltzmann method was later decoupled from the lattice gas entirely when Qian, d’Humieres, and Lallemand [5] proposed a simplified BGK collision operator. By simply choosing a local equilibrium distribution (the now ubiquitous second order polynomial equilibrium distribution) they simplified the algorithm and made it significantly more accessible. They were also able to remove some non-Galilean invariant artifacts of Boolean lattice gas systems, such as a velocity dependent viscosity [1, 6] and pressure, even if these deviations became most notable in parameter ranges (i.e. at high Mach numbers) that were still not accessible to these lattice Boltzmann methods.
@@
The simplicity of this BGK collision operator, more than anything else, influenced the explosive rise of lattice Boltzmann methods over their ancestral lattice gas methods. Certain systems, however, such as the Brownian motion of colloids considered by Ladd et al. [7], require fluctuations to be simulated. These fluctuating systems were, at the time, limited to being simulated by lattice gases. To allow these simulations to benefit from the advances of lattice Boltzmann methods, Ladd developed a fluctuating lattice Boltzmann method. The basic idea was to implement Landau and Lifshitz’ [8] fluctuating stress tensor into the collision operator leading to a Langevin approach [9] for the fluctuating lattice Boltzmann method. This fluctuating lattice Boltzmann method recovered the fluctuating hydrodynamic equations [10]. However, utilizing the fluctuating stress tensor alone limited the recovery of the correct fluctuations to those with a wavelength on the order of the system size, while the total order of fluctuations was suppressed by about a factor of two.
@@
Adhikari et al. discovered that to recover the smaller scale fluctuations, a full set of fluctuating modes, rather than just those associated with hydrodynamic quantities, was needed [11]. This method is restricted to ideal gas systems. For non-ideal systems additional corrections are required [12]. Since the Langevin approach is based on continuous densities, the large fluctuations at low densities will lead to the formation of non-physical negative densities [13, sec. 5.3].
@@
When considering methods for simulating fluctuating systems, however, it is much more natural to think in terms of discrete systems since fluctuations fundamentally arise because of the discreteness of nature [14]. This inspired Chopard et al. to instead develop novel lattice gas algorithms instead of adding fluctuations artificially to the lattice Boltzmann algorithm [15, 16, sec. 5.7]. They extended the occupation numbers of the lattice gas from Boolean values to integer values, a concept that was first introduced by Chen et al. the year prior with their Digital Physics approach [17]. This allowed for the lattice gas system of Chopard et al. to recover a Poisson distribution in global equilibrium. Secondly, Chopard et al. implemented a sampling algorithm, which was originally theorized by Boghosian et al. [18], that distributed particles randomly across the system velocities based on a local equilibrium ensemble. For the diffusive case, a multinomial distribution over the occupation numbers themselves was utilized for the local equilibrium ensemble, which was sufficient to follow the conservation of mass required by the system [16, Sec. 5.7].
@@
In extending the method to hydrodynamic systems, however, Chopard et al. did not utilize an explicit local equilibrium ensemble due to the complexity involved in deriving this distribution while accounting for momentum conservation. Instead, a multinomial distribution similar to that of the diffusive case was used, centered around the second order polynomial equilibrium distribution. This on its own would violate momentum conservation, so a random factor that redistributed a single particle at a time until momentum conservation was upheld was also added [15]. Both of these methods were, for higher densities, well approximated by Gaussian distributions, which caused their algorithmic complexity to be approximately O(1) with respect to the system density, matching the complexity of the lattice Boltzmann algorithm.
@@
Blommel and Wagner created another integer based lattice gas algorithm to simulate hydrodynamic systems [19]. Their algorithm utilized a series of random two-particle collisions to redistribute particle velocities while still maintaining momentum conservation with collision rules explicitly derived by imposing detailed balance on the system. This allowed Blommel and Wagner to recover Poisson distributed fluctuations in global equilibrium similar to Chopard et al..
@@
When analyzing their algorithm, the global equilibrium average distribution was found to correspond not to the polynomial lattice Boltzmann equilibrium distribution, but to the entropic lattice Boltzmann equilibrium distribution. This equilibrium distribution had already been derived by Ansumali et al. [20] previously by minimizing an H-functional. The polynomial equilibrium distribution utilized by Chopard et al. [15] for their hydrodynamic multiparticle lattice gas, therefore, was actually incompatible with a microscopic lattice gas implementation.
@@
Since the number of individual two-particle collisions needed to achieve a state close to local equilibrium scales with the square of the number of particles at each lattice site, Blommel and Wagner’s algorithm proved much slower than the fluctuating lattice Boltzmann algorithm even for moderate local densities of a few hundred particles per lattice site. This lead Seekins and Wagner [21] to accelerate Blommel’s algorithm. A sampling method similar to that of Chopard et al. was implemented to reduce the number of random numbers needed, leading to increased efficiency.
@@
Using detailed balance to derive a local equilibrium distribution, Seekins and Wagner recovered an identical multinomial distribution to that of Chopard et al. [16, Sec. 5.7]. Although this meant that the diffusive algorithm created by Seekins and Wagner was essentially identical to that of Chopard et al., the derivation of the equilibrium distribution via detailed balance was novel, and we show here that it can be used to derive the proper equilibrium distribution for hydrodynamic systems. The efficiency of the algorithm was comparable to a diffusive fluctuating lattice Boltzmann algorithm, along with having a complexity at high densities of approximately O(1) with respect to density [21, 22].
@@
In this paper, we extend our prior work for diffusive systems to hydrodynamic systems by deriving an appropriate local equilibrium distribution. For simplicity, we limit the system to one dimension, leaving an extension to higher dimensions to future work. The paper is structured as follows: in Section II of the paper, we will cover the sampling lattice gas algorithm in detail and derive the local equilibrium distribution and provide an analysis of the Poisson distribution as a global equilibrium distribution for the integer lattice gas. Section III will discuss the Boltzmann limit of the lattice gas and its uniqueness. We show that there isn’t a unique Boltzmann limit, although for many practical situations the BGK collision operator can be a good approximation for the Boltzmann limit. Section IV covers the results of two test cases used to study the non-equilibrium dynamics of the system. Mostly we demonstrate excellent agreement between the averaged lattice gas and the equivalent lattice Boltzmann approach. However, in certain situations there is clear evidence of lattice gas behavior beyond the Boltzmann limit. Finally, Section V will compare the efficiency of our sampling lattice gas to that of both Blommel and Wagner and an entropic lattice Boltzmann algorithm showing that the new implementation of the integer lattice gas shows significant speed up to make it comparable in efficiency with the entropic lattice Boltzmann method.
@@ II.1 Sampling Lattice Gas Basics
A lattice gas consists of a set of discrete cells situated on a lattice, each containing a discrete number of velocity states $v_{i}$ for particles to occupy [23]. The velocities are usually identical for every cell, and are defined such that if $x$ is a position on the lattice, then for any velocity $v_{i}$, $x+v_{i}$ is also a position on the lattice, assuming a timestep of $\Delta t=1$. These velocities are also typically restricted in number such that the particle displacement over a single timestep is small. The particles are accounted for by the occupation numbers $n_{i}(x,t)$, denoting the particles that move from lattice site $x-v_{i}$ at time t1t-1 to lattice site $x$ at time $t$. These particles are then redistributed among the occupation numbers, observing the applicable conservation rules, and streamed. This two-step process represents a discrete timestep for the system. Thus, the $n_{i}(x,t)$ evolve according to,
@@
which is known as the lattice gas evolution equation. Here $\{n_{j}\}_{g}$ represents the set of all $n_{i}$ of the system at the current time, and $\{n_{j}\}_{l}$ represents the set of $n_{i}$ at a single lattice site at the current time, with the subscripts $l$ and $g$ denoting between the two. In this evolution equation, the collision operator $\Xi_{i}$ is a stochastic operator. The explicit nature of the $n_{i}$ is dependent on the lattice gas system observed. For Boolean lattice gases, for example, these $n_{i}$ would be Boolean in nature, leading to $\Xi_{i}$ to be calculable explicitly within a lookup table due to the limited nature of the probability space for single-occupation velocity states [24, 1]. In our integer lattice gas method, the $n_{i}$ values become integer based, which leads to a collision operator $\Xi_{i}$ that can take on integer values.
In the ideal gas systems considered here the collision operator will only act on the local distribution of occupation numbers, $\{n_{j}\}_{l}$, which consists only of the local occupation numbers at position $x$ and $t$. In this nomenclature we exclude the implicit dependence on $x$ and $t$ to make the equations more readable. However, in principle the collision operator can depend on the global distribution $\{n_{j}\}_{g}$, as would be the case for non-ideal lattice gas systems. An example of that would be Rothmann and Keller’s algorithm [25] that is now known as the color gradient method. This method depends on neighboring site average density gradients.
The sampling collision operator takes inspiration from the BGK collision operator for the Boltzmann equation. The Boltzmann equation is an evolution equation for a distribution function $f_{i}(x,t)$, and the BGK collision operator is an approximation that simply moves the system towards it’s local equilibrium distribution $f_{i}^{0}(x,t)$ [26] where the local equilibrium distribution depends on the locally conserved quantities. This BGK lattice Boltzmann algorithm will be discussed in more detail in upcoming sections.
In a lattice gas system, the system is described by $\{n_{i}\}_{g}$. Unlike in the lattice Boltzmann case, the local equilibrium state for the $n_{i}$ is not given by an equilibrium value, but rather by a probability distribution around some average value, known as the local equilibrium ensemble. This local equilibrium ensemble, given a local pre-collision set of occupation numbers $\{n_{i}\}_{l}$, gives a probability $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ for each post-collision set $\{\hat{n}_{i}\}_{l}$. A post-collision set of occupation numbers can be then be sampled from the local equilibrium ensemble in a process discussed in detail in Seekins and Wagner’s paper [21]. The collision operator for the sampling lattice gas algorithm thus becomes the difference between the pre-collision and post-collision occupation numbers, explicitly defined as,
@@
However, this representation of the collision operator does not account for the fact that the system may not relax fully to local equilibrium in one timestep. To simulate this under-relaxation in a system of discrete particles, we can represent the fraction of particles that should undergo collision in that timestep as $\omega$. We create a subset of our occupation numbers $\{n^{\omega}_{i}\}_{l}$ to undergo collision, with the probability that any single particle is a member of the subset being given by $\omega$. Therefore, each specific value of $n_{i}^{\omega}$ has a probability associated with it given by the binomial distribution,
@@
All that remains to fully define the collision operator is to derive the local equilibrium ensemble explicitly.
@@ II.2 Derivation of the Local Equilibrium Ensemble
In the hydrodynamic case considered here there are two conserved quantities: mass and momentum. Since momentum is conserved, the probability distribution for the $n_{i}$ will be correlated. However, by transforming our representation of the system from being expressed as an set of individual occupation numbers $\{n_{i}\}_{l}$ into being expressed by the hydrodynamic moments we may be able to obtain an uncorrelated probability distribution. These moments are physical quantities within the simulation, with a total number given by the total number of system velocities $Q$. From some fundamental particle collision process like that of Blommel and Wagner [19] we deduce the local equilibrium probability distributions for the non-conserved moments. We then use these local equilibrium ensembles to sample the post-collision moments as random variables. This is the extent of how far we are able to describe our algorithm in all generality.
To know what moment representation is applicable to our system, we must focus our attention on a specific example. As this paper serves as a proof of concept for our sampling algorithm, we focus on the simplest possible system that recovers hydrodynamic effects, a one-dimensional system with three total velocity vectors, leaving the extension of the algorithm to more complex systems, including higher dimensional systems, to future work.
@@
Thus, we can redefine $P^{0}(\{\hat{n}_{i}\}_{l};\{n_{i}\}_{l})$ in terms of $N$, $J$, and $\pi$ as $P^{0}(\hat{N},\hat{J},\hat{\pi};N,J,\pi)$, or the probability of finding a system in a state with moments $\hat{N}$, $\hat{J}$, and $\hat{\pi}$ after a collision given a pre-collision state with moments $N$, $J$, and $\pi$. We can then simplify this representation, as our conservation rules state that $\hat{N}=N$ and $\hat{J}=J$. Furthermore, because we are assuming that the system is in local equilibrium when described by this distribution, we know that $\hat{\pi}$ does not depend on the pre-collision $\pi$ value. The probability distribution for our ensemble can then be written as $P^{0}(\pi;N,J)$, which is the equilibrium probability of finding the system in a post-collision state with a $\pi$ value of $\pi$, given a pre-collision system with conserved moments $N$ and $J$, and an arbitrary initial value for $\pi$.
We now derive the particular probability distribution that recovers the integer lattice gas results obtained by Blommel and Wagner. In local equilibrium, this distribution must be invariant under collisions. By using Blommel and Wagner’s algorithm as a basis, we restrict the possible collisions of our system to only involving pairs of particles. These collisions conserve momentum, and their equilibrium ensemble must be stationary over each collision. This restricts us from altering states where two selected particles are moving in the same direction, such as $\{1,1\}$, where changing the velocity of the particles would violate momentum conservation in 1D systems. Since we are using momentum space to track the collision, any collisions that do not alter the $\pi$ value of the system have no effect, an example being $\{1,-1\}\to\{-1,1\}$. There are only four two-particle collisions for a D1Q3 system that alter the $\pi$ value: two that increase the $\pi$ value by two and two that decrease it by two,
[…]
The probabilities of each of these possible collisions occurring is given by the probability of picking a pair of particles with the proper set of velocity indices $\{i,j\}$, which is written as $p_{\{i,j\}}$, multiplied by the probability of the collision occurring from velocity state $\{i,j\}$ to velocity state $\{k,l\}$, written as $P_{\{i,j\}\to\{k,l\}}$. Thus, we can write $T_{\pi\to\pi+2}$, which is the probability of the system transitioning from a state with a $\pi$ value of $\pi$ to one with a $\pi$ value of $\pi+2$, as,
@@
and $T_{\pi+2\to\pi}$, which is the probability of the system transitioning from a state with a $\pi$ value of $\pi+2$ to one with a $\pi$ value of $\pi$, as,
@@
where $\delta_{ij}$ is the Kronecker delta, and is incorporated to avoid double counting particles should the indices $i$ and $j$ be the same. The $P_{\{i,j\}\to\{k,l\}}$ values were defined explicitly by Blommel and Wagner such that at zero velocity the typical lattice Boltzmann equilibrium values are recovered on average [19]. For the relevant collisions in Eqs. (13) and (14), these values are given by,
@@
It is not obvious that the probability $P_{\{i,j\}\to\{k,l\}}$ should necessarily be independent of the locally conserved quantities. However, as Blommel and Wagner’s method makes this assumption, we make it as well to adapt their algorithm. A true test of the validity of this approach could be done by looking at the coarse-graining of a molecular dynamics simulation, similar to the work done by Parsa et al. [27].
@@
We can write,
$T_{\pi\to\pi+2}P^{0}(\pi;N,J)=T_{\pi+2\to\pi}P^{0}(\pi+2;N,J),$ (24)
using the fact that $P^{0}(\pi;N,J)$ must be invariant under a single collision. Eq. (24) can then be rewritten using Eqs. (22) and (23),
@@
fully describes the local equilibrium ensemble. If we sample random $\hat{\pi}$ values out of this local equilibrium ensemble, our collision operator in moment space becomes,